OpenStudy (anonymous):

3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (anonymous):

-3

hero (hero):

y = (3-3x)/-6 y = (11+7x)/5 y = y -66-42x = 15 - 15x 27x = -81 x = -3 y = -2

OpenStudy (anonymous):

He is just interested in the answer!!!!

hero (hero):

I like posting the solution anyway

OpenStudy (anonymous):

He doesn't want to learn he just wants to have the easy way out!